Imagine that you work in a World Bank office that is in charge of making recommendations to increase the growth of the economy in Nigeria. According to the information you have available, the population of Nigeria is growing at an annual rate of 5%, capital depreciates at 3% annually, the savings rate is 10% and Nigeria's economy requires an additional unit of capital to produce an additional unit of product. Your direct boss tells you that in the office they use the Harrod-Domar model.

After a conversation with your boss you convince him to use the Solow model (without technological change) to better understand the balance that Nigeria's economy could achieve in the long term. The information you have allows you to know that Nigeria's production function is such that y = 10,000 k, where y is the per-capita GDP and k is the per-capita capital. In equilibrium, what is the value of y and k for Nigeria?
Answer #1

In case of solow model, at steady state

s*y = (n+d)*k

y = 10,000√k

.1y = (.05 + .03)k

So

.1*10,000√k = .08k

1000/.08 = √k

√k* = 12,500

k* = (12,500)2

so eqm k = 156,250,000

Eqm y = 10,000*√k

= 125,000,000